OpenStudy (anonymous):

write the equation in the point slope form (5,1);m=1/3 (-2,3);m=-2

OpenStudy (anonymous):

y - y1 = m(x - x1)

OpenStudy (anonymous):

@littleRob wanna give it a try?

OpenStudy (anonymous):

3y=x-2 y=-2x-1

OpenStudy (anonymous):

y=mx+b m is 1/3 plug the value (5, 1) to find b and then u have the equation in point slope form do the same thing with the second part as well if u have any trouble u can send me a message because i dont always pay attention to all replies or someone else can help u I hope it helps